OpenStudy (anonymous):

What is the least common denominator of the expression below? x^2 + 6 / x^2 + 4x - 12 + 7x / x^2 + 7x + 6 a. x^2 + 4x + 4 b. (x + 6) c. (x + 6)(x - 2) (x + 1) d. (x + 6)(x - 2)(x + 6)(x + 1) help please.

jimthompson5910 (jim_thompson5910):

I'll get you started \[\Large \frac{x^2+6}{x^2+4x-12} + \frac{7x}{x^2+7x+6}\] \[\Large \frac{x^2+6}{(x-2)(x+6)} + \frac{7x}{x^2+7x+6}\] \[\Large \frac{x^2+6}{(x-2)(x+6)} + \frac{7x}{(x+6)(x+1)}\] and I'll let you finish up

OpenStudy (anonymous):

C

jimthompson5910 (jim_thompson5910):

Again, the idea is to step 1) factor all the denominators step 2) multiply all of the unique factors in the denominators to get the LCD

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

:D